A contractor orders 8 cubic yards of premixed cement, all of which is to be used to pour a patio that will be 4 inches thick. If the length of the patio is specified to be twice the width, what will be the patio dimensions?
Answer by checkley77(12844)

8yd^3=8*3*3*3=216 ft^3
Volume=l*w*d (length, width, depth)
216=2w*w*1/3
216=2w^2/3
2w*2=216*3
2w^2=648
w^2=648/2
w^2=324
w=18 ft. is the width.
2*18=36 ft. is the length.
Proof
18*36*1/3=216
648/3=216
216=216